OpenStudy (anonymous):

Find x if (2/7)^2x = (7/2)^3x + 5.

OpenStudy (anonymous):

need help

OpenStudy (anonymous):

is the + 5 an exponent?

OpenStudy (anonymous):

oooooooooooooooh

OpenStudy (anonymous):

i thought it was \[(\frac{2}{7})^{2x}=(\frac{7}{2})^{3x}+5\]

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

^3x+5

OpenStudy (anonymous):

but instead it is \[(\frac{2}{7})^{2x}=(\frac{7}{2})^{3x+5}\]?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

that we can do! use logs

OpenStudy (anonymous):

\[2x\ln(\frac{2}{7})=(3x+5)\ln(\frac{7}{2})\]

OpenStudy (anonymous):

is step one. the algebra from here on in

OpenStudy (anonymous):

what is In?

OpenStudy (anonymous):

\[\ln(\frac{2}{7})\times 2x=\ln(\frac{7}{2})\times 3x+\ln(\frac{7}{2})\times 5\] etc.

OpenStudy (anonymous):

what is ln?

OpenStudy (anonymous):

well if you have never seen it this problem will not make sense unless you want to replace ln by log

OpenStudy (anonymous):

do you know what log is?

OpenStudy (anonymous):

i don't understand what is In?

OpenStudy (anonymous):

do you know log? as in \[f(x)=\log(x)\]?

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

i am asking.

OpenStudy (anonymous):

I don't know

OpenStudy (zarkon):

\[\left(\frac{2}{7}\right)^{2x}=\left(\frac{7}{2}\right)^{3x+5}\] \[\left(\frac{7}{2}\right)^{-2x}=\left(\frac{7}{2}\right)^{3x+5}\] \[\Rightarrow -2x=3x+5\Rightarrow -5=5x\Rightarrow x=-1\]

OpenStudy (anonymous):

yes it is

OpenStudy (anonymous):

thanks